OpenStudy (anonymous):

How do you solve this? Determine which values of the variable must be excluded from the domain 5/2x+1.

OpenStudy (phi):

you are not allowed to divide by zero if there is an x value that makes the denominator zero, you must exclude that x value

OpenStudy (phi):

in other words, if 2x + 1 = 0 that would be bad. what x value would make 2x+1 = 0 ?

OpenStudy (phi):

you can use algebra to find the "bad" x value. 2x + 1 =0 add -1 to both sides: 2x + 1 -1 = 0 -1 or 2x = -1 divide both sides by 2 2x/2 = -1/2 or x = -1/2

OpenStudy (anonymous):

i just noticed i typed in the wrong numbers! ugh! the write problem is 3x+7/4x-3

OpenStudy (anonymous):

the same rules apply, right?

OpenStudy (anonymous):

Yes. But be careful with higher order polynomaials. When you factorize, things may cancel.

OpenStudy (phi):

When you factorize, things may cancel. True, but for the simplified expression to be the same as the original it must have the same domain. It must exclude any values the "cancelled"

OpenStudy (phi):

the same rules apply, right? Yes. 4x-3 cannot be 0 can you solve for the x value where 4x - 3 = 0 ?

OpenStudy (anonymous):

would the answer be 3/4?

OpenStudy (anonymous):

well x= 3/4?

OpenStudy (phi):

x= 3/4 must be excluded from the domain

OpenStudy (anonymous):

Ah yes. With the factorizing, be careful to still consider the "cancelled" terms Guess what I said was misleading

OpenStudy (anonymous):

thank you Phi and Xavier :)

OpenStudy (phi):

If you get any square roots, another rule is no square roots of a negative number.

OpenStudy (anonymous):

awesome, thanks.